# Avoiding Catastrophe in Online Learning by Asking for Help

### v4

@@ 3 Model
$|\underbrace{\mu_{t}(x,\pi^{m}(x))}_{\text{Taking the right action}}-\underbrace{\mu_{t}(x,\pi^{m}(x^{\prime}))}_{\text{Using what you learned in $x^{\prime}$}}|\ \ \ \leq\underbrace{L||x-x^{\prime}||}_{\text{Similarity between $x$ and $x^{\prime}$}}$
Borrowing knowledge from similar experiences seems fundamental to learning and is well-understood in the psychology literature (Esser et al., 2023) and education literature (Hajian, 2019).
Crucially, our input space can be any feature embedding of the agent’s situation, not just its physical positioning. Our algorithms will not require knowledge of the feature embedding and do not need to know $L$, so it suffices that there exists some feature embedding which satisfies local generalization. The agent does not even need to know which embedding it is. Finally, local generalization implies the more familiar Lipschitz continuity for an optimal mentor (Proposition E.1).
Multiplicative objective and regret.
@@
If $\mu_{t}(x_{t},y_{t})\in[0,1]$ is the chance of avoiding catastrophe on time step $t$ (conditioned on no prior catastrophe), then $\prod_{t=1}^{T}\mu_{t}(x_{t},y_{t})$ is the agent’s overall chance of avoiding catastrophe by the chain rule of probability. For given $\boldsymbol{x},\boldsymbol{y},\boldsymbol{\mu},\pi^{m}$, the agent’s regret is
$R_{T}(\boldsymbol{x},\boldsymbol{y},\boldsymbol{\mu},\pi^{m})=\prod_{t=1}^{T}\mu_{t}^{m}(x_{t})-\prod_{t=1}^{T}\mu_{t}(x_{t},y_{t})$
@@
We write $R_{T}=R_{T}(\boldsymbol{x},\boldsymbol{y},\boldsymbol{\mu},\pi^{m})$ for brevity. We desire subconstant worst-case regret: the total (not average) expected regret should go to 0 for any $\boldsymbol{\mu}$ and $\pi^{m}$; formally, $\lim_{T\to\infty}\sup_{\boldsymbol{\mu},\pi^{m}}\operatorname*{\mathbb{E}}[R_{T}]=0$.
The interpretation of a bound on $\operatorname*{\mathbb{E}}[R_{T}]$ depends on the quality of the mentor. For example, subconstant regret becomes trivial if $\lim_{T\to\infty}\operatorname*{\mathbb{E}}\big{[}\prod_{t=1}^{T}\mu_{t}^{m}(x_{t})\big{]}=0$. However, we think that high-stakes AI applications should ensure the presence of a mentor who is almost always safe, i.e., $\operatorname*{\mathbb{E}}\big{[}\prod_{t=1}^{T}\mu_{t}^{m}(x_{t})\big{]}\approx 1$. If no such mentor exists for some application, perhaps it is better to avoid the application altogether. Also, our regret bounds include rates of convergence, so even if the mentor policy will eventually cause catastrophe, we can still bound how quickly the agent becomes unsafe.
VC and Littlestone dimensions.
@@
VC dimension (Vapnik & Chervonenkis, 1971) and Littlestone dimension (Littlestone, 1988) are standard measures of learning difficulty which capture the ability of a hypothesis class (in our case, a policy class) to realize arbitrary combinations of labels (in our case, actions). We omit the precise dimensions since we only utilize these concepts via existing results. See Shalev-Shwartz & Ben-David (2014) for a comprehensive overview.
Misc.
@@
The diameter of a set $X\subseteq\mathcal{X}$ is defined by $\operatorname*{diam}(X)=\max_{x,x^{\prime}\in X}||x-x^{\prime}||$. All logarithms and exponents are base $e$ unless otherwise noted.
4 Avoiding catastrophe is impossible in general
We begin by showing that in general, any algorithm with sublinear mentor queries has arbitrary poor regret in the worst-case, even when inputs are i.i.d. on $[0,1]$. The result also holds even if the algorithm knows $L$ and $\boldsymbol{x}$ ahead of time. The proof can be found in Appendix A.
@@ 4 Avoiding catastrophe is impossible in general
The worst-case expected regret of any algorithm with sublinear queries goes to 1 as $T$ goes to infinity. Formally, $\lim_{T\to\infty}\ \,\sup_{\boldsymbol{\mu},\pi^{m}}\,\operatorname*{\mathbb{E}}[R_{T}]=1$.
@@ 4.1 Intuition
We partition $\mathcal{X}$ into equally-sized sections that are “independent” in the sense that querying an input in section $i$ gives you no information about section $j$. The number of sections is determined by a function $f:\mathbb{N}\to\mathbb{N}$ that we will choose. If $|Q_{T}|\in o(f(T))$, most of these sections will never contain a query. When the agent sees an input in a section not containing a query, it essentially has to guess, meaning it will be wrong a constant fraction of the time. We then choose a payoff function (which is the same for all time steps) which makes the wrong guesses as costly as possible, subject to the local generalization constraint. Figure 2 fleshes out this idea.
Picking $\boldsymbol{f(T)}$.
A natural idea is to try $f(T)=T$, but this doesn’t quite work: even if the agent chooses wrong on every time step, the minimum payoff is still at least $1-\frac{L}{2T}$, and $\lim_{T\to\infty}\prod_{t=1}^{T}\left(1-\frac{L}{2T}\right)=\lim_{T\to\infty}\left(1-\frac{L}{2T}\right)^{T}=e^{-L/2}$. In order for the regret to approach 1, we need $f(T)$ to be asymptotically between $|Q_{T}|$ and $T$ (such $f$ must exist since $|Q_{T}|\leq g(T)\in o(T)$). This leads to the following bound: $\prod_{t=1}^{T}\mu(x_{t},y_{t})\leq\big{(}1-\frac{L}{\Theta(f(T))}\big{)}^{\Theta(T)}$. When $f(T)\in o(T)$, the right hand side converges to 0, while $\prod_{t=1}^{T}\mu^{m}(x_{t})=1$. In words, the agent is nearly guaranteed to cause catastrophe, despite the existence of a policy which is guaranteed to avoid catastrophe.
$x$$\mu$  $\mu(x,0)$$\mu(x,1)$$\cfrac{L}{2f(T)}$$\frac{1}{f(T)}$$\frac{2}{f(T)}$$\frac{3}{f(T)}$$\frac{f(T)-1}{f(T)}$$\dots$1001111
Figure 2: An illustration of the construction we use to prove Theorem 4.1 (not to scale). The horizontal axis indicates the input $x\in[0,1]$ and the vertical axis indicates the payoff $\mu(x,y)\in[0,1]$. The solid line represents $\mu(x,0)$ and the dotted line represents $\mu(x,1)$. In each section, one of the actions has the optimal payoff of 1, and the other action has the worst possible payoff allowed by $L$, reaching a minimum of $1-\frac{L}{2f(T)}$. Crucially, both actions result in a payoff of 1 at the boundaries between sections: this allows us to “reset” for the next section. As a result, we can freely toggle the optimal action for each section independently.
VC dimension.
The class of mentor policies induced by our construction has VC dimension $f(T)$; considered over all possible values of $T$, this implies infinite VC (and Littlestone) dimension. This is necessary given our positive results in Section 5.
@@ 4.2 Formal definition of construction
Let $\mathcal{X}=[0,1]$ and $\mathcal{D}_{t}=U$ for each $t\in[T]$. Assume that $L\leq 1$; this will simplify the math and only makes the problem easier for the agent. We define a family of payoff functions parameterized by a function $f:\mathbb{N}\to\mathbb{N}$ and a bit string $\boldsymbol{a}=(a_{1},a_{2},\dots,a_{f(T)})\in\{0,1\}^{f(T)}$. The bit $a_{j}$ will denote the optimal action in section $j$. Note that $f(T)\geq 1$ and since we defined $\mathbb{N}$ to exclude 0.
[…]
@@
We will use this payoff function for all time steps: $\mu_{t}=\mu_{f,\boldsymbol{a}}$ for all $t\in[T]$. Let $\pi^{m}$ be any policy which is optimal for $\mu_{f,\boldsymbol{a}}$. Note that there is a unique optimal action for each $x_{t}$, since each $x_{t}$ belongs to exactly one $X_{j}$; formally, $\pi^{m}(x_{t})=a_{j(x_{t})}$.
@@ 5 Avoiding catastrophe given finite VC or Littlestone dimension
Theorem 4.1 shows that avoiding catastrophe is impossible in general, which is also true in online learning. What if we restrict ourselves to settings where standard online learning is possible? Specifically, we assume that $\pi^{m}$ belongs to a policy class $\Pi$ where either (1) $\Pi$ has finite VC dimension $d$ and $\boldsymbol{x}$ is $\sigma$-smooth or (2) $\Pi$ has finite Littlestone dimension $d$ [^10] Recall from Section 1.3 that standard online learning becomes tractable under either of these assumptions.
This section presents a simple algorithm which guarantees subconstant regret and sublinear queries under either of those assumptions. Our algorithm needs to know $\Pi$, as is standard in online learning. The algorithm does not need to know $\sigma$ (in the smooth case) or $L$, and can handle an unbounded input space (the number of queries simply scales with the maximum distance between observed inputs).
For simplicity, we initially prove our result for $\mathcal{Y}=\{0,1\}$ (the proof can be found in Appendix B). Appendix C extends our result to many actions using the standard “one versus rest” reduction. [^11] For each action $y$, we learn a binary classifier which predicts whether $\pi^{m}(x)=y$. If every binary classifier is correct, we can correctly determine $\pi^{m}(x)$. See, e.g., Chapter 29 of Shalev-Shwartz & Ben-David (2014).
@@ 5.1 Intuition behind the algorithm
Algorithm 1 has two simple components: (1) run a modified version of the Hedge algorithm for online learning, but (2) ask for help for unfamiliar inputs (specifically, when the input is very different from any queried input with the same action under the proposed policy). Hedge ensures that the number of time steps where the agent’s action doesn’t match the mentor’s is small, and asking for help for unfamiliar inputs ensures that when we do make a mistake, the cost isn’t too high. This algorithmic structure seems quite natural: mostly follow a baseline strategy, but ask for help when out-of-distribution.
Agnostic to feature embedding.
The algorithm does not need to know how inputs are represented in $\mathbb{R}^{n}$, since the only nontrivial operation it performs on the input space is computing a nearest neighbor distance (line 21 of Algorithm 1). This could be modeled as an out-of-distribution detector score, for which many methods are available (see e.g., Yang et al. (2024)). This agnosticism reduces the burden on the local generalization assumption: it suffices that there exists some feature embedding satisfying this assumption.
Hedge.
Hedge (Freund & Schapire, 1997) is a standard online learning algorithm which ensures sublinear regret when the number of hypotheses (in our case, the number of policies in $\Pi$) is finite. [^12] See Chapter 5 of Slivkins et al. (2019) and Chapter 21 of Shalev-Shwartz & Ben-David (2014) for modern introductions to Hedge. We would prefer not to assume that $\Pi$ is finite. Luckily, any policy $\Pi$ can be approximated within $\varepsilon$ when either (1) $\Pi$ has finite VC dimension and $\boldsymbol{x}$ is $\sigma$-smooth or (2) $\Pi$ has finite Littlestone dimension. Thus we can run Hedge on this approximative policy class instead.
@@
1:function AvoidCatastrophe($T\in\mathbb{N},\>\varepsilon\in\mathbb{R}_{>0},\>d\in\mathbb{N},\,$ policy class $\Pi$)
2:     if $\Pi$ has VC dimension $d$ then
3:         $\tilde{\Pi}\leftarrow$ any smooth $\varepsilon$-cover of $\Pi$ of size at most $(41/\varepsilon)^{d}$triangleright See Definition 5.4
4:     else
5:         $\tilde{\Pi}\leftarrow$ any adversarial cover of size at most $(eT/d)^{d}$triangleright See Definition 5.5
6:     $X\leftarrow\emptyset$
7:     $w(\pi)\leftarrow 1$ for all $\pi\in\tilde{\Pi}$
8:     $p\leftarrow 1/\sqrt{\varepsilon T}$
9:     $\eta\leftarrow\max\big{(}\sqrt{\frac{p\log|\tilde{\Pi}|}{2T}},\>\frac{p^{2}}{\sqrt{2}}\big{)}$
10:     for $t$
from
$1$
to
$T$ do $\triangleright$ Run one step of Hedge, which selects policy $\pi_{t}$

### v5

@@ 3 Model
$\big{|}\!\!\underbrace{\mu_{t}(x,\pi^{m}(x))}_{\text{Taking the right action}}-\!\underbrace{\mu_{t}(x,\pi^{m}(x^{\prime}))}_{\text{Using what you learned in $x^{\prime}$}}\!\!\!\!\!\!\!\big{|}\,\leq\>\underbrace{L||x-x^{\prime}||}_{\text{Input similarity}}$
This ability seems fundamental to intelligence and is well-understood in psychology (e.g., Esser et al., 2023) and education (e.g., Hajian, 2019). Note that the input space $\mathcal{X}\subseteq\mathbb{R}^{n}$ can be any encoding of the agent’s situation, not just its physical positioning. See Section 5.2 for further discussion.
All suprema over $\boldsymbol{\mu},\pi^{m}$ pairs are assumed to be restricted to $\boldsymbol{\mu},\pi^{m}$ pairs which satisfy local generalization.
@@
Regret. If $\mu_{t}(x_{t},y_{t})\in[0,1]$ is the chance of avoiding catastrophe at time $t$ (conditioned on no prior catastrophe), then by the chain rule of probability, $\prod_{t=1}^{T}\mu_{t}(x_{t},y_{t})$ is the agent’s overall chance of avoiding catastrophe. For given $\boldsymbol{x},\boldsymbol{y},\boldsymbol{\mu},\pi^{m}$, the agent’s multiplicative regret [^6] One could also define the multiplicative regret as $R_{T}^{\prime}=\prod_{t=1}^{T}\mu_{t}^{m}(x_{t})-\prod_{t=1}^{T}\mu_{t}(x_{t},y_{t})$, but our definition is actually stricter: $\lim_{T\to\infty}R_{T}^{\times}\to 0$ implies $\lim_{T\to\infty}R_{T}^{\prime}\to 0$, while the reverse is not true. In particular, $\lim_{T\to\infty}R_{T}^{\prime}\to 0$ is trivial if $\lim_{T\to\infty}\prod_{t=1}^{T}\mu_{t}^{m}(x_{t})\to 0$. is
$R_{T}^{\times}(\boldsymbol{x},\boldsymbol{y},\boldsymbol{\mu},\pi^{m})=\log\prod_{t=1}^{T}\mu_{t}^{m}(x_{t})-\log\prod_{t=1}^{T}\mu_{t}(x_{t},y_{t})$
when all payoffs are strictly positive. To handle the case where some payoffs are zero, we assume the existence of $\mu_{0}^{m}>0$ such that $\mu^{m}(x_{t})\geq\mu_{0}^{m}$ always. Thus only the agent’s payoffs can be zero, so we can safely define $R_{T}^{\times}(\boldsymbol{x},\boldsymbol{y},\boldsymbol{\mu},\pi^{m})=\infty$ whenever $\mu_{t}(x_{t},y_{t})=0$ for some $t\in[T]$. We write $R_{T}^{\times}=R_{T}^{\times}(\boldsymbol{x},\boldsymbol{y},\boldsymbol{\mu},\pi^{m})$ for brevity.
The assumption of $\mu^{m}(x_{t})\geq\mu_{0}^{m}>0$ means that the mentor cannot be abysmal. In fact, we argue that high-stakes AI applications should employ a mentor who is almost always safe, i.e., $\mu_{0}^{m}\approx 1$. If no such mentor exists for some application, perhaps that application should be avoided altogether.
We also define the agent’s additive regret as
$R_{T}^{+}(\boldsymbol{x},\boldsymbol{y},\boldsymbol{\mu},\pi^{m})=\sum_{t=1}^{T}\mu_{t}^{m}(x_{t})-\sum_{t=1}^{T}\mu_{t}(x_{t},y_{t})$
@@
and similarly write $R_{T}^{+}=R_{T}^{+}(\boldsymbol{x},\boldsymbol{y},\boldsymbol{\mu},\pi^{m})$ for brevity. For both objectives, we desire subconstant worst-case regret: the total (not average) expected regret should go to 0 for any $\boldsymbol{\mu}$ and $\pi^{m}$. Formally, we want $\lim_{T\to\infty}\sup_{\boldsymbol{\mu},\pi^{m}}\operatorname*{\mathbb{E}}[R_{T}^{\times}]=0$ and $\lim_{T\to\infty}\sup_{\boldsymbol{\mu},\pi^{m}}\operatorname*{\mathbb{E}}[R_{T}^{+}]=0$.
@@
VC and Littlestone dimensions. VC dimension (Vapnik & Chervonenkis, 1971) and Littlestone dimension (Littlestone, 1988) are standard measures of learning difficulty which capture the ability of a hypothesis class (in our case, a policy class) to realize arbitrary combinations of labels (in our case, actions). We omit the precise definitions since we only utilize these concepts via existing results. See Shalev-Shwartz & Ben-David (2014) for a comprehensive overview.
@@
Misc. The diameter of a set $S\subseteq\mathcal{X}$ is defined by $\operatorname*{diam}(S)=\max_{x,x^{\prime}\in S}||x-x^{\prime}||$. All logarithms and exponents are base $e$ unless otherwise noted.
input $x$payoff $\mu$   $\mu(x,0)$$\mu(x,1)$$\cfrac{L}{2f(T)}$$\frac{1}{f(T)}$$\frac{2}{f(T)}$$\frac{3}{f(T)}$$\frac{f(T)-1}{f(T)}$$\dots$1001111
Figure 1: An illustration of the construction we use to prove Theorem 4.1 (not to scale). The horizontal axis indicates the input $x\in[0,1]=\mathcal{X}$ and the vertical axis indicates the payoff $\mu(x,y)\in[0,1]$. The solid line represents $\mu(x,0)$ and the dotted line represents $\mu(x,1)$. In each section, one of the actions has the optimal payoff of 1, and the other action has the worst possible payoff allowed by $L$, reaching a minimum of $1-\frac{L}{2f(T)}$. Crucially, both actions result in a payoff of 1 at the boundaries between sections: this allows us to “reset” for the next section. As a result, we can freely toggle the optimal action for each section independently.
4 Avoiding catastrophe with sublinear queries is impossible in general
We first show that in general, any algorithm with sublinear mentor queries has unbounded regret in the worst-case, even when inputs are i.i.d. on $[0,1]$ and $\boldsymbol{\mu}$ does not vary over time. The formal proofs are deferred to Appendix A, but we provide intuition and define the construction here.
@@ 4 Avoiding catastrophe with sublinear queries is impossible in general
Any algorithm with sublinear queries has unbounded worst-case regret (both multiplicative and additive) as $T\to\infty$. Specifically,
$\sup_{\boldsymbol{\mu},\pi^{m}}\operatorname*{\mathbb{E}}[R_{T}^{\times}],\,\sup_{\boldsymbol{\mu},\pi^{m}}\operatorname*{\mathbb{E}}[R_{T}^{+}]\in\Omega\left(\!L\sqrt{\frac{T}{\sup_{\boldsymbol{\mu},\pi^{m}}\operatorname*{\mathbb{E}}[|Q_{T}|]\!+\!1}}\right)$
Intuitively, the regret decreases as the number of queries increases. However, as long as the number of queries remains sublinear in $T$, the regret is unbounded as $T\to\infty$.
We also have the following corollary of Theorem 4.1:
Corollary 4.1.1.
Even when $\mu_{t}^{m}(x)=1$ for all $t$ and $x$, any algorithm with sublinear queries satisfies
$\lim_{T\to\infty}\,\sup_{\boldsymbol{\mu},\pi^{m}}\,\operatorname*{\mathbb{E}}\left[\prod_{t=1}^{T}\mu_{t}(x_{t},y_{t})\right]=0$
In other words, even if the mentor never causes catastrophe, any algorithm with sublinear queries causes catastrophe with probability 1 as $T\to\infty$ in the worst case.
@@ 4.1 Intuition
We partition $\mathcal{X}$ into equally-sized sections that are “independent” in the sense that querying an input in section $i$ provides no information about section $j$. There will be $f(T)$ sections, where $f$ is a function that we will choose. If $|Q_{T}|\in o(f(T))$, most of these sections will never contain a query. When the agent sees an input in a section not containing a query, it essentially must guess, meaning it will be wrong about half the time. We then choose a payoff function (which is the same for all time steps) which makes the wrong guesses as costly as possible, subject to the local generalization constraint. Figure 1 fleshes out this idea.
The choice of $f$ is crucial. One idea is $f(T)=T$. If the agent is wrong about half the time, and the average payoff for wrong actions is $1-\frac{L}{4T}$, we can estimate the regret as
$R_{T}^{\times}\,=$ $\ \log\prod_{t=1}^{T}\mu_{t}^{m}(x_{t})-\log\prod_{t=1}^{T}\mu_{t}(x_{t},y_{t})$
$\approx$ $\ \log 1-\log\left(1-\frac{L}{4T}\right)^{T/2}$
$=$ $\ -\frac{T}{2}\log\left(1-\frac{L}{4T}\right)$
$\approx$ $\ \frac{T}{2}\cdot\frac{L}{4T}$
$=$ $\ \frac{L}{8}$
Thus $f(T)=T$ can at best give us a constant lower bound on regret. Instead, we choose $f$ such that $|Q_{T}|\in o(f(T))$ and $f(T)\in o(T)$. Specifically, we choose $f(T)=\max(\sqrt{\sup_{\boldsymbol{\mu},\pi^{m}}\operatorname*{\mathbb{E}}[|Q_{T}|]T},1)$. Most sections still will not contain a query, so the agent is still wrong about half the time, but the payoff for wrong actions is worse. Then
$R_{T}^{\times}\,\approx$ $\ \log 1-\log\left(1-\frac{L}{4f(T)}\right)^{T/2}$
$\approx$ $\ \frac{LT}{8f(T)}$
$\in$ $\ \Omega\left(L\sqrt{\frac{T}{\sup_{\boldsymbol{\mu},\pi^{m}}\operatorname*{\mathbb{E}}[|Q_{T}|]+1}}\right)$
which produces the bound in Theorem 4.1.
VC dimension. The class of mentor policies in our construction has VC dimension $f(T)$; across all possible values of $T$, this implies infinite VC (and Littlestone) dimension. We know that this is necessary given our positive results.
@@ 4.2 Formal definition of construction
Let $\mathcal{X}=[0,1]$ and $\mathcal{D}_{t}=U(\mathcal{X})$ for each $t\in[T]$, where $U(\mathcal{X})$ is the uniform distribution on $\mathcal{X}$. Assume that $L\leq 1$; this will simplify the math and only makes the problem easier for the agent. We define a family of payoff functions parameterized by a function $f:\mathbb{N}\to\mathbb{N}$ and a bit string $\boldsymbol{a}=(a_{1},a_{2},\dots,a_{f(T)})\in\{0,1\}^{f(T)}$. The bit $a_{j}$ will denote the optimal action in section $j$. Note that $f(T)\geq 1$.
For each $j\in[f(T)]$, we refer to $X_{j}=\left[\frac{j-1}{f(T)},\frac{j}{f(T)}\right]$ as the $j$th section. Let $m_{j}=\frac{j-0.5}{f(T)}$ be the midpoint of $X_{j}$. Assume that each $x_{t}$ belongs to exactly one $X_{j}$ (this happens with probability 1, so this assumption does not affect the expected regret). Let $j(x)$ denote the index of the section containing input $x$. Then $\mu_{f,\boldsymbol{a}}$ is defined by
@@
We use this payoff function for all time steps: $\mu_{t}=\mu_{f,\boldsymbol{a}}$ for all $t\in[T]$. Let $\pi^{m}$ be any optimal policy for $\mu_{f,\boldsymbol{a}}$. Note that there is a unique optimal action for each $x_{t}$, since each $x_{t}$ belongs to exactly one $X_{j}$; formally, $\pi^{m}(x_{t})=a_{j(x_{t})}$.
@@ 5 Avoiding catastrophe given finite VC or Littlestone dimension
Theorem 4.1 shows that avoiding catastrophe is impossible in general. What if we restrict ourselves to settings where standard online learning is possible? Specifically, we assume that $\pi^{m}$ belongs to a policy class $\Pi$ where either (1) $\Pi$ has finite VC dimension $d$ and $\boldsymbol{x}$ is $\sigma$-smooth or (2) $\Pi$ has finite Littlestone dimension $d$ [^7] Recall from Section 1.3 that standard online learning becomes tractable under either of these assumptions. This section presents a simple algorithm which guarantees subconstant regret (both multiplicative and additive) and sublinear queries under either of those assumptions. Formal proofs are deferred to Appendix B but we provide intuition and a proof sketch here.
@@ 5.1 Intuition behind the algorithm
Algorithm 1 has two simple components: (1) run a modified version of the Hedge algorithm for online learning, but (2) ask for help for unfamiliar inputs (specifically, when the input is very different from any queried input with the same action under the proposed policy). Hedge ensures that the number of mistakes (i.e., the number of time steps where the agent’s action doesn’t match the mentor’s) is small, and asking for help for unfamiliar inputs ensures that when we do make a mistake, the cost isn’t too high. This algorithmic structure seems quite natural: mostly follow a baseline strategy, but ask for help when out-of-distribution.
Hedge. Hedge (Freund & Schapire, 1997) is a standard online learning algorithm which ensures sublinear regret when the number of hypotheses (in our case, the number of policies in $\Pi$) is finite. [^8] Chapter 5 of Slivkins et al. (2019) and Chapter 21 of Shalev-Shwartz & Ben-David (2014) give modern introductions to Hedge. We would prefer not to assume that $\Pi$ is finite. Luckily, any policy in $\Pi$ can be approximated within $\varepsilon$ when either (1) $\Pi$ has finite VC dimension and $\boldsymbol{x}$ is $\sigma$-smooth or (2) $\Pi$ has finite Littlestone dimension. Thus we can run Hedge on this approximative policy class instead.
@@
5.2 Local generalization
Local generalization is vital: this is what allows us to detect when an input is unfamiliar. Crucially, our algorithm does not need to know how inputs are encoded in $\mathbb{R}^{n}$ and does not need to know $L$: it only needs to be able to compute the nearest-neighbor distance $\min_{(x,y)\in S:y=\pi_{t}(x_{t})}||x_{t}-x||$. Thus we only need to assume that there exists some encoding satisfying local generalization.
To elaborate, recall the example that a 3 mm spot and a 3.1 mm spot on X-rays likely have similar risk levels (assuming similar density, location, etc.). If the risk level abruptly increases for any spot over 3 mm, then local generalization may not hold for a naive encoding which treats size as a single dimension. However, a more nuanced encoding would recognize that these two situations – a 3 mm vs 3.1 mm spot – are in fact not similar. Constructing a suitable encoding may be challenging, but we do not require the agent to have explicit access to such an encoding: the agent only needs a nearest-neighbor distance oracle.
  Inputs: $T\in\mathbb{N},\>\varepsilon\in\mathbb{R}_{>0},\>d\in\mathbb{N},\,$ policy class $\Pi$
  if $\Pi$ has VC dimension $d$ then
     $\tilde{\Pi}\leftarrow$ any smooth $\varepsilon$-cover of $\Pi$ of size at most $(41/\varepsilon)^{d}$  (see Definition 5.3)
  else
     $\tilde{\Pi}\leftarrow$ any adversarial cover of size at most $(eT/d)^{d}$  (see Definition 5.4)
  $S\leftarrow\emptyset$
  $w(\pi)\leftarrow 1$ for all $\pi\in\tilde{\Pi}$
  $p\leftarrow 1/\sqrt{\varepsilon T}$
  $\eta\leftarrow\max\big{(}\sqrt{\frac{p\log|\tilde{\Pi}|}{2T}},\>\frac{p^{2}}{\sqrt{2}}\big{)}$
  for $t$ from $1$ to $T$ do
     Run one step of Hedge, which selects policy $\pi_{t}$